A magnetic field of 3.2T is perpendicular to a square coil of 12 turns.The length of each side of the coil is 9cm.Find the magnetic flux through the coil if the magnetic field makes an angle of 30° with the normal to the plane of the coil.

1
Expert's answer
2022-04-11T12:03:57-0400

The magnetic flux

"\Phi=NBA\cos\theta\\
=12*3.2*0.09^2*\cos30^\circ=0.27\:\rm Wb"